sove for x sin3y = 4 siny * sin(x+y) *sin(x-y) where y is not equal to npi

 
sove for x
sin3y = 4 siny * sin(x+y) *sin(x-y) where y is not equal to npi

Grade:11

2 Answers

Nicho priyatham
7 years ago
sin3y = 4 siny * sin(x+y) *sin(x-y)
 sin(x+y) *sin(x-y)=sin2 x-sin2 y
so sin3y= 4siny(sin2 x-sin2 y)
3siny-4sin3y=4sinysin2x-4sin3y
so sin2x=3/4
x=npie+ pie/3
PLZ approve if usefull
Incase of any doubts drop it in answer box
 
 
raj
7 years ago
hello, 
sin3y = 4 siny * sin(x+y) *sin(x-y)
 sin(x+y) *sin(x-y)=sin2 x-sin2 y
so sin3y= 4siny(sin2 x-sin2 y)
3siny-4sin3y=4sinysin2x-4sin3y
so sin2x=3/4
x=npie+ pie/3